OpenStudy (anonymous):

What is the solution to the equation? -7 ( x - 4 ) + 8x = 12 + 14

OpenStudy (anonymous):

x = -2 -7(x-4) +8x =26 -7x+28+8x = 26 x+28=26 x=26-28

OpenStudy (anonymous):

Thank you !!!